For a sample size of 7 and success rate of 5%, find the probability of:

a)6 successes

b)1 success

c)at least 5 successes

d)more than 5 successes

e)at most 4 successes
Answer #1

X ~ Bin ( n , p)

Where n = 6 , p = 0.05

Binomial probability distribution is

P(X) = nCx * px * ( 1 - p)n-x

a)

P(X = 6) = 7C6 * 0.056 * ( 1 - 0.05)1

= 0.0000

b)

P(X = 1) = 7C1 * 0.051 * ( 1 - 0.05)6

= 0.2573

c)

P(C >= 5) = P(X =5) + P(X = 6)

= 7C5 * 0.055 * ( 1 - 0.05)2 +7C6 * 0.056 * ( 1 - 0.05)1

= 0.0000

d)

P(X > 5) = P(X = 6)

= 7C6 * 0.056 * ( 1 - 0.05)1

= 0.0000

e)

P(X <= 4) = 1 - P(X > = 5)

= 1 - 0000

= 1